# CL-Polyp: A Contrastive Learning-Enhanced Network for Accurate Polyp Segmentation

### v1

@@ 3 Methods
In this paper, we propose that CL-Polyp consists of two branches: the segmentation and the contrastive learning branches. The main branch is the polyp segmentation branch, and the addition of contrastive learning is the auxiliary. The implementation details are provided below.
Figure 3: A ResNet50 Encoder extracts the feature map of an image and then enters it into two branches. One branch is used for the segmentation task. The other branch is compared with the features obtained by the momentum updating Encoder with positive and negative samples. The MLP of the contrast learning branch is composed of the Linear layer, BN layer, ReLU layer, and Linear layer. The grey box part represents the inference process.
@@ 3.1.1 Selection of positive and negative samples
The training by contrastive learning is similar to the training method of the multi-task model structure, which uses a mixed loss function to train the network structure. Contrastive learning focuses on constructing positive and negative sample pairs for data. In this paper, according to the classification method of data in TGANet[4], polyps are divided into one or more according to the number of polyps and the size. According to the number of polyps, the data can be divided into one or many; According to the size of polyps, they can be classified into small, medium, and large. Combining the number and size categories, the total can be divided into six categories, as shown in Fig.1.
$N$ and $S$ respectively, polyp number and size of the collection, polyp size, and the number of data sets described as polyps set $C=N\times S=\{(n,s)|n\in N,s\in S\}$. Training set data $D_{tr}=\{(x,c)|c\in C\}$. Through this division, it is very convenient to select positive and negative samples.
@@
There are two ways to select positive samples $x_{p}^{i}$ similar to the original image, one is data augmentation of image and the other is from $c=\{(n,s)|n=\hat{n},s=\hat{s}\}$ choose one of the data as the positive example. In our experiments, we choose one kind of data augmentation of the image as the positive sample. The number and size of polyps in the positive sample are the same as in this data. Negative selection and the image size and the number of samples of different samples, the samples of negative samples $x_{n}^{i}$ to choose $\{(n,s)|n\neq\hat{n},s\neq\hat{s}\}$.
@@ 3.1.2 Detailed structure
As shown in Fig.3,the Encoder comprises a ResNet and an MLP. Specifically, ResNet50 is used. The MLP consists of a Linear layer, BN layer, ReLU activation function, and Linear layer. The image passes through the last layer block of ResNet, after which it is divided into two branches, one branch is used for contrastive learning, and the other branch enters the main segmentation branch. After average pooling, the branch features of contrastive learning obtain 2048 dimensional vector features and, finally, 2048 dimensional vector features after MLP. Specifically, $x^{i}$ extracted the visual features $h$ of the polyp image through the Encoder, $x_{p}^{i}$, $x_{n}^{i}$ respectively by a momentum encoder of the same structure through the structure respectively after obtaining $h_{p}$, $h_{n}$. The parameters of the momentum encoder are updated utilizing momentum in the following way.
@@
Then we use contrastive learning loss function to optimize these three feature vectors. The overall training process is presented in Algorithm.1. For the inference procedure, as shown in the gray box in Fig.3, there are no positive and negative samples and the process of momentum Encoder.
@@ 3.1.3 Contrastive Learning Loss Function
The contrastive learning method ensures that the features of a polyp image $X^{i}$ are similar to the positive sample $X_{p}^{i}$, and For image features, the distance between image features and positive features is always closer than the distance between features of negative samples and image features. The loss function of contrastive learning is introduced to optimize $h$, $h_{p}$, and $h_{n}$ to reduce the distance between $h$ and $h_{p}$. At the same time, the distance between $h$, and each negative sample $h_{n}$ is increased. The triplet loss function is adopted, which is proposed in Face-Net[43] and the specific loss function can be described as
@@
Where $\alpha$ is the interval between positive and negative samples. $N$ is all instances of the training set. In the actual algorithm, one image data corresponds to one positive sample and $K$ negative samples. In my experiment, $K$ is set to 4.
3.1.4 Choices for data augmentation
Figure 4: Graph a presents the overall structure of the segmentation model which is the inference procedure of Fig.3. We use ResNet50 as Encoder, in Decoder we use our proposed MASPP module and CA module. Graph b shows the detailed structure of the MASPP module.
Learning data invariance is crucial for contrastive learning[44]. Proper selection of data augmentation is essential for understanding data invariance, which directly impacts the model’s learning ability. In polyp segmentation tasks, common data augmentation methods include image resizing and cropping, rotation, horizontal flipping, color transformation, and grayscale change. However, we selected random cropping, resizing, rotation, horizontal flipping, Gaussian noise, and normalization as our data augmentation techniques.
@@ 3.2.1 Modified ASPP module
The ASPP module proposed in DeppLabV3+, ASPP module can obtain multi-scale information through multiple dilated convolution modules with different proportions. In ASPP, channel concatenates, and 1x1 convolution layers are used to fuse multi-scale information after multiple dilated convolutions. Still, the fusion of multi-scale information can only partially be carried out by using channel concatenate. Therefore, we propose a module called Modified Atrous Spatial Pyramid Pooling (MASPP module) which can better fuse high-level multi-scale features.
Fig.4b shows the detailed MASPP module. The high-level feature information $f_{5}$, and then two branch structures, the first branch structure is a ASPP module, after multiple dilated convolution modules, the channel concatenate and 1x1 convolution are simply fused. The other branch structure is to maintain the original feature structure, the features passed through 3x3 convolution layers, BN layers, and ReLU activation layers. After that, an SE module, proposed in SENet[29], is used to fuse these two branch features. The SE module includes Squeeze & Excitation. In the Squeeze module,the first average pooling turn the feature $f_{5}\in\mathbb{R}^{C\times H\times W}$ into a shape of $C\times 1\times 1$. Then through a Linear layer, the ReLU activation layer, Linear layer becomes a shape of $H\times W\times 1\times 1$ and then the original data and the activated vector are multiplied, which can better fuse multiple features. This can be expressed as follows.
@@
3.2.2 Using the CA module
In Deeplabv3+ model, skip connection improves the boundary information of the second layer features in Decoder. While using only one layer of features is great, this ignores features from other layers. We propose the CA module to better fuse the features into the Decoder. The second layer feature $f_{2}\in\mathbb{R}^{C\times H\times W}$ , $l$ represents the second number of layers of the feature, after two 1 x 1 convolution will output two features $A_{i},B_{i}$, and the upsample feature $z_{i}$, concrete can be expressed as $z_{i}=A_{i}+\rm Conv(\rm cat(B_{i},z_{i}))$. A 4-fold upsampling operation and an Outconv then obtain the final prediction result.
@@ 3.2.3 Segmentation branch loss function
To reduce overfitting, we adopt the BCEDice loss used in nnU-Net[46]. It consists of Dice loss, and BCE loss. Dice loss can be widely applied to medical image tasks and is written as follows.
@@
Where $i$ denotes the index of all pixels in the probability distribution map. $P_{i}$ is the probability that the $i$-th pixel belongs to the segmented region, and $G_{i}$ is the true value of the $i$-th pixel. Finally, the loss function of the segmentation part consists of two parts: BCEloss and Diceloss.
@@ 3.3 Hybrid loss function
The overall loss function consists of two parts: the loss function of contrastive learning and the loss function of segmentation.
@@
4 Experiment
4.1 Datasets and Experiment Scenarios
@@
The Kvasir-SEG dataset consists of 1000 images. The images in the dataset are irregular in size, ranging from 332x487 to 1920x1072. The size of the polyp ranged from 0.79 % to 62.13 %, and the shape and size of the polyp were irregular. So it’s the most challenging. The CVC-ClinicDB dataset includes 612 images obtained from 29 polyp videos. The image size is 384 x 288. The size varies from 0.34% to 45.88%. The CVC-ColonDB dataset includes 380 images obtained from 16 colonoscopy videos with an image size of 574 x 500. Polyp size from 0.30% to 63.15 %. The CVC300 dataset s a cross-domain dataset that contains 60 polyp images. The ETIS dataset includes 196 images obtained from 16 enteroscopy videos. The image size is 1225 x 996. It has a larger resolution compared to other images.
@@
Experiment Scenario III: We conduct ablation experiments on the Kvasi-SEG dataset. To verify the effectiveness of the various modules proposed in our method, the training set and test set are divided in the same way as that of Experiment Scenario I. Ablation experiments on the Kvaisr-SEG dataset can reveal the effectiveness of each part of our method. In addition, we conducted experiments to verify the impact of different data augmentations on the model.
4.2 Experiment settings
The experiments use the PyTorch framework on Ubuntu22.04, using python version 3.9. The hardware device, using a single NVIDIA 1080TI GPU with 11GB memory, accelerates our segmentation model’s training with a mini-batch size of 4. Using the Adam optimizer, the model used an initial learning rate of 1e-4 to train our model. Learning Rate Decay Strategy Cosine learning rate decay strategy is chosen. Besides, the input images are resized to 384 x 384. Due to reasons with our training method, we trained 300 epochs for both Experiments Scenario I, Experiment Scenario II, and Experiment Scenario III.
@@ 4.3 Evaluation metrics
The most common medical image segmentation task uses the Dice coefficient and IoU as evaluation indicators. Most models use this to evaluate how good the model is. We used Dice and IoU as evaluation indexes in Experimental Scenario I, II, and III. The dice coefficient is a standard measure to compare the predicted label with the actual value. It can be defined as:
@@
IoU is another standard segmentation evaluation measure used to calculate the overlap between the predicted and true values. According to the following formula definition:
@@
TP, FP, TN, and FN represent True Positive, False Positive, True Negative, and False Negative respectively. In addition, we also compared three indexes, namely precision, recall, and F2 score in Experiment Scenario I.
@@
4.4 Experiment result
4.4.1 The result of Experiment Scenario I
@@
Table 1 shows the results on the Kvasir-SEG dataset. The comparison of these models all achieved good performance, with Dice above 0.90 and IoU above 0.83. Some methods failed to reach the highest accuracy in the paper, because Experiment Scenario I was only conducted on the Kvasir-SEG dataset. The PraNet method achieved a Dice of 0.894 and IoU of 0.830. The average Dice of TransFuse-S/L was 0.90 and IoU was 0.84. GLFRNet (TMI’22) had the highest Dice of 0.911 and IoU of 0.853. The best result of our proposed method is that Dice reaches 0.918 and IoU reaches 0.864, which is higher than TGANet which uses the same partitioning method as us. Both metrics of our methods Precision and F2 are also the highest.
Table 1: Experiment Scenario I: Results of ten medical image segmentation methods on the Kvasir-SEG dataset.
@@
To show our model intuitively, we provide results of our model on Kvasir-SEG test images and compare testing results with six advanced models. As shown in Fig.5, the proposed method can obtain better outcomes for the input image in the first four rows, regardless of the overall polyp or the edge details. Our approach can distinguish multiple polyps clearly for the figure’s last three rows of pictures, while other models are challenging to determine. Especially for the polyps with fuzzy boundaries in the last row, our method can still effectively segment the polyps.
4.4.2 The results of Experiment Scenario II
@@
Table 2: Results of 10 medical image segmentation methods on Kvasir-SEG and CVC-ClinicDB datasets by the training method of PraNet.
@@
Table 3: Results of ten medical image segmentation methods on the CVC-ColonDB, CVC-300 and ETIS datasets by the training method of PraNet.‘-’ denotes that the corresponding value is not reported.
@@
Figure 5: Visual comparison between the proposed method and six state-of-the-art methods.
4.4.3 The results of Experiment Scenario III
@@
Table 4: The comparison of our CL-Polyp with different modules on Kvasir-SEG datasets.
@@
Table 5: The comparison of our CL-Polyp with different data augmentation on Kvasir-SEG datasets.
@@
Data enhancement methods are essential for learning data invariance by contrast learning, so we explored the influence of different data enhancement methods on our methods. Table 5 shows the experimental results, where Base data enhancement refers to random resize and crop, and the range of crop is (0.8, 1.0), random rotation, vertical flip, and normalization. Based on base data enhancement, we added gaussian blur, to gray, brightness contrast, and their random combination and other contents. The ablation results show that only adding the gaussian blur method to the Base data enhancement method can achieve the best results.

### v2

@@ 3 Methods
In this paper, we propose that CL-Polyp consists of two branches: the segmentation and the contrastive learning branches. The main branch performs polyp segmentation, while contrastive learning serves as an auxiliary task. The implementation details are provided below.
Figure 3: Architecture of the ResNet50-based segmentation model with contrastive learning: Feature maps extracted by the ResNet50 encoder are fed into a segmentation branch and a contrastive learning branch. The contrastive branch compares features with those from a momentum-updated encoder using positive and negative samples. The multi-layer perceptron (MLP) in the contrastive branch consists of linear, batch normalization (BN), and ReLU layers.
@@ 3.1.1 Selection of positive and negative samples
The training by contrastive learning is similar to the training method of the multi-task model structure, which uses a mixed loss function to train the network structure. Contrastive learning focuses on constructing positive and negative sample pairs for data. In this paper, according to the classification method of data in TGANet[4], polyps are divided into one or more according to the number of polyps and the size. Based on polyp count, data can be divided into single or multiple polyps; based on size, they can be classified as small, medium, or large. Combining the number and size categories, the total can be divided into six categories, as shown in Fig.1.
$N$ and $S$ respectively, polyp number and size of the collection, polyp size, and the number of data sets described as polyps set $C=N\times S=\{(n,s)|n\in N,s\in S\}$. Training set data $D_{tr}=\{(x,c)|c\in C\}$. This categorization facilitates the selection of positive and negative samples.
@@
There are two ways to select positive samples $x_{p}^{i}$ similar to the original image, one is data augmentation of image and the other is from $c=\{(n,s)|n=\hat{n},s=\hat{s}\}$ choose one of the data as the positive example. In our experiments, we use data augmentation of the image to create the positive sample. The positive sample has the same number and size of polyps as the original data. Negative selection and the image size and the number of samples of different samples, the samples of negative samples $x_{n}^{i}$ to choose $\{(n,s)|n\neq\hat{n},s\neq\hat{s}\}$.
@@ 3.1.2 Detailed structure
As shown in Fig.3,the Encoder comprises a ResNet and an MLP. Specifically, ResNet50 is used. The MLP consists of a Linear layer, BN layer, ReLU activation function, and Linear layer. The image passes through ResNet’s last layer block, then splits into two branches: one for contrastive learning and the other for the main segmentation branch. After average pooling, the contrastive learning branch features are processed by the MLP to obtain a 2048 -dimensional vector. Specifically, the Encoder extracts visual features $h$ from the polyp image $x^{i}$, Then, $x_{p}^{i}$ and $x_{n}^{i}$ are passed through a momentum encoder with the same structure to obtain $h_{p}$ and $h_{n}$, respectively. The parameters of the momentum encoder are updated utilizing momentum in the following way.
@@
Then, we use the contrastive learning loss function to optimize these three feature vectors. The overall training process is presented in Algorithm.1. For the inference procedure, as shown in the gray box in Fig.3, there are no positive and negative samples and the process of momentum Encoder.
@@ 3.1.3 Contrastive Learning Loss Function
The contrastive learning method ensures that the features of a polyp image $X^{i}$ are similar to the positive sample $X_{p}^{i}$, such that the distance between image features and positive sample features is always less than the distance between negative sample features and image features. The loss function of contrastive learning is introduced to optimize $h$, $h_{p}$, and $h_{n}$ to reduce the distance between $h$ and $h_{p}$. Concurrently, the distance between $h$ and each negative sample $h_{n}$ is increased. The triplet loss function is adopted, which is proposed in Face-Net[53] and the specific loss function can be described as
@@
Where $\alpha$ is the interval between positive and negative samples. $N$ is all instances of the training set. In practice, each image corresponds to one positive sample and $K$ negative samples. In our experiments, $K$ is set to 4.
3.1.4 Data Augmentation Techniques
Figure 4: Segmentation model architecture: (a) Overall structure with ResNet50 encoder, MASPP, and CA modules in the decoder. (b) Detailed structure of the MASPP module.
Learning data invariance is crucial for contrastive learning[54]. Proper selection of data augmentation is essential for understanding data invariance, which directly impacts the model’s learning ability. In polyp segmentation tasks, common data augmentation methods include image resizing and cropping, rotation, horizontal flipping, color transformation, and grayscale change. We selected random cropping, resizing, rotation, horizontal flipping, Gaussian noise, and normalization as our data augmentation techniques.
@@ 3.2.1 Modified ASPP module
The ASPP module proposed in DeppLabV3+, ASPP module can obtain multi-scale information through multiple dilated convolution modules with different proportions. In ASPP, channel concatenates, and 1x1 convolution layers are used to fuse multi-scale information after multiple dilated convolutions. Still, the fusion of multi-scale information can only partially be carried out by using channel concatenate. Therefore, we propose the Modified Atrous Spatial Pyramid Pooling (MASPP) module to improve the fusion of high-level multi-scale features.
Fig.4b shows the detailed MASPP module. The high-level feature information $f_{5}$, and then two branch structures, the first branch structure is a ASPP module, after multiple dilated convolution modules, the channel concatenate and 1x1 convolution are simply fused. The other branch maintains the original feature structure by passing the features through 3x3 convolution layers, BN layers, and ReLU activation layers. After that, an SE module, proposed in SENet[39], is used to fuse these two branch features. The SE module includes Squeeze & Excitation. In the Squeeze module,the first average pooling turn the feature $f_{5}\in\mathbb{R}^{C\times H\times W}$ into a shape of $C\times 1\times 1$. Then through a Linear layer, the ReLU activation layer, Linear layer becomes a shape of $H\times W\times 1\times 1$ and then the original data and the activated vector are multiplied, which can better fuse multiple features. This can be expressed as follows.
@@
3.2.2 CA Module
In the DeepLabV3+ model, skip connections enhance boundary information using second -layer Decoder features. However, this ignores other layers. We propose the CA module to improve feature fusion in the Decoder. The second layer feature $f_{2}\in\mathbb{R}^{C\times H\times W}$ , $l$ represents the second number of layers of the feature, after two 1 x 1 convolution will output two features $A_{i},B_{i}$, and the upsample feature $z_{i}$, concrete can be expressed as $z_{i}=A_{i}+\rm Conv(\rm cat(B_{i},z_{i}))$. A 4-fold upsampling operation and an Outconv then obtain the final prediction result.
@@ 3.2.3 Segmentation branch loss function
To reduce overfitting, we adopt the BCEDice loss used in nnU-Net[56]. It consists of Dice loss and BCE loss. Dice loss can be widely applied to medical image tasks and is written as follows.
@@
Where $i$ denotes the index of all pixels in the probability distribution map. $P_{i}$ is the probability that the $i$-th pixel belongs to the segmented region, and $G_{i}$ is the true value of the $i$-th pixel. Finally, the segmentation loss function combines BCE loss and Dice loss.
@@ 3.3 Hybrid loss function
The overall loss function combines the contrastive learning loss and the segmentation loss.
@@
4 Experiments and Results
4.1 Datasets and Scenarios
@@
The Kvasir-SEG dataset consists of 1000 images. The images in the dataset are irregular in size, ranging from 332x487 to 1920x1072. The size of the polyp ranged from 0.79 % to 62.13 %, and the shape and size of the polyp were irregular. So it’s the most challenging. The CVC-ClinicDB dataset includes 612 images obtained from 29 polyp videos. The image size is 384 x 288. The size varies from 0.34% to 45.88%. The CVC-ColonDB dataset includes 380 images obtained from 16 colonoscopy videos with an image size of 574 x 500. Polyp size from 0.30% to 63.15 %. The CVC-300 dataset is a cross-domain dataset containing 60 polyp images. The ETIS dataset includes 196 images obtained from 16 enteroscopy videos. The image size is 1225 x 996. It has a larger resolution compared to other images.
@@
Experiment Scenario III: We conduct ablation experiments on the Kvasi-SEG dataset. To verify the effectiveness of the various modules proposed in our method, the training set and test set are divided in the same way as that of Experiment Scenario I. Ablation experiments on the Kvaisr-SEG dataset can reveal the effectiveness of each part of our method. Additionally, we experimented to verify the impact of different data augmentations on the model.
4.2 Experimental Setup
The experiments use the PyTorch framework on Ubuntu22.04, using python version 3.9. The hardware device, using a single NVIDIA 1080TI GPU with 11GB memory, accelerates our segmentation model’s training with a mini-batch size of 4. Using the Adam optimizer, the model used an initial learning rate of 1e-4 to train our model. Learning Rate Decay Strategy Cosine learning rate decay strategy is chosen. Besides, the input images are resized to 384 x 384. Due to our training method, we trained for 300 epochs in Experiment Scenarios I, II, and III.
@@ 4.3 Evaluation Metrics
Medical image segmentation commonly uses the Dice coefficient and IoU as evaluation metrics. We used Dice and IoU as evaluation metrics in Experimental Scenarios I, II, and III. The Dice coefficient is a standard measure for comparing the predicted segmentation with the ground truth and can be defined as:
@@
IoU, another standard segmentation evaluation metric, calculates the overlap between the predicted and ground truth values, as defined by the following formula:
@@
TP, FP, TN, and FN represent True Positive, False Positive, True Negative, and False Negative respectively. Additionally, we compared three metrics—precision, recall, and the F2 score —in Experiment Scenario I.
@@
4.4 Experimental Results
4.4.1 Results of Experiment Scenario I
@@
Table 1 shows the results on the Kvasir-SEG dataset. The comparison of these models all achieved good performance, with Dice above 0.90 and IoU above 0.83. Some methods failed to reach the highest accuracy in the paper, because Experiment Scenario I was only conducted on the Kvasir-SEG dataset. The PraNet method achieved a Dice of 0.894 and IoU of 0.830. The average Dice of TransFuse-S/L was 0.90 and IoU was 0.84. GLFRNet (TMI’22) had the highest Dice of 0.911 and IoU of 0.853. The best result of our proposed method is that Dice reaches 0.918 and IoU reaches 0.864, which is higher than TGANet which uses the same partitioning method as us. Our method also achieved the highest precision and F2 score.
Table 1: Segmentation performance of ten methods on the Kvasir-SEG dataset in Experiment Scenario I: Comparative results.
@@
To show our model intuitively, we provide results of our model on Kvasir-SEG test images and compare testing results with six advanced models. As shown in Fig.5, the proposed method can obtain better outcomes for the input image in the first four rows, regardless of the overall polyp or the edge details. Our approach can distinguish multiple polyps clearly for the figure’s last three rows of pictures, while other models are challenging to determine. Even for polyps with fuzzy boundaries in the last row, our method effectively segments them.
4.4.2 Results of Experiment Scenario II
@@
Table 2: Segmentation performance of ten methods on the Kvasir-SEG and CVC-ClinicDB datasets using the PraNet training method: Comparative analysis of results.
@@
Table 3: Segmentation performance of ten methods on the CVC-ColonDB, CVC-300, and ETIS datasets using the PraNet training method: Comparative analysis of results. Missing values are indicated by ‘-’.
@@
Figure 5: Qualitative comparison of segmentation results: Proposed method versus six state-of-the-art methods on representative images.
4.4.3 Results of Experiment Scenario III
@@
Table 4: Performance comparison of CL-Polyp with different modules on the Kvasir-SEG dataset: Ablation study results.
@@
Table 5: Impact of different data augmentation techniques on CL-Polyp performance using the Kvasir-SEG dataset: Comparative results.
@@
Data enhancement methods are essential for learning data invariance by contrast learning, so we explored the influence of different data enhancement methods on our methods. Table 5 shows the experimental results, where Base data enhancement refers to random resize and crop, and the range of crop is (0.8, 1.0), random rotation, vertical flip, and normalization. Based on base data enhancement, we added gaussian blur, to gray, brightness contrast, and their random combination and other contents. The ablation results show that adding only Gaussian blur to the Base data augmentation yields the best results.